Sharing Walt’s Dream

An Unofficial Guide To Disneyland

By Roger B. Jones, Jr. September 1995

This guide is dedicated to my mother, who taught me that being a child is not reserved for the young, and to my wife,

who shows me that I don’t ever want to grow up.

This guide may be distributed freely so long as it is not modified in any way and all terms stated in the “Legal Stuff”

section are complied with.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994,1995 Roger B. Jones, Jr. All Rights Reserved.

Introduction

Man has a dream, and that’s a start He follows his dream with mind and heart And when it becomes a reality It’s a dream come true for you and me

– There’s A Great Big Beautiful Tomorrow

That quote, in four lines, tells the story of Disneyland. Disneyland began with the dream of one man, Walt Disney. He chased his dream despite the objections of many who “knew better,” who knew that it was “a crazy idea.” When, after years of work, his dream became a reality, it became “The Happiest Place on Earth.”

This guide is meant to help you enjoy a trip into Disneyland. The guide is given to all, to help everyone share Walt’s dream.

To all who come to this happy place: Welcome. Disneyland is your land. Here age relives fond memories of the past… and here youth may savor the challenge and promise of the future. Disneyland is dedicated to the ideals, the dreams, and the hard facts that have created America… with the hope that it will be a source of joy and inspiration to all the world.

– Walt Disney’s dedication of Disneyland, July 17, 1955

About The Guide

Getting The Guide

You’ve obviously figured this out to some extent, since you’re reading this now. However, in case you got an older version, the latest version of the guide can be obtained by any of the following methods:

    FTP: ftp://ftp.io.com:/pub/usr/kazrak/Disney/Disneyland

Email: send mail to rjones@us.oracle.com with the subject line “SEND TEXT” for an ASCII version, “SEND PS” for PostScript, “SEND DVI” for the TEX DVI file, or “SEND HTML” for the HTML version. Due to changes in how the guide is made, I can no longer offer it in Microsoft Word format.

Mail: send $3 to the address below, along with a note giving your address. This money is just to cover expenses of printing and postage.

In addition, I’ve set up a mailing list to distribute information when new versions come out. Send email to rjones@us.oracle.com with the subject line “JOIN GUIDE LIST” to be added to the list.

Author’s Notes

This book started with another book, The Very Unofficial Guide to WDW by Todd McCartney. Here was a book, freely available over the Internet, that had just as much information as the travel guides costing $12 and up! I decided that a similar guide was needed for Disneyland.

I have been to the Magic Kingdom four times, each time for one entire day (opening to closing). For additional information (since four times is actually a rather low number for Disneyland visits), I have relied on my wife, who has been somewhere over fifteen times, and on information picked up from others. I have made every attempt to ensure that the information in this book is accurate; however, since Disneyland is constantly changing, some information may be out of date. For comments on the book, or to correct an error, please contact me at:

Postal Mail:  Roger B. Jones, Jr.

244 N. Ellsworth Ave. #10 San Mateo, CA 94401

Electronic Mail: rjones@us.oracle.com; kazrak@io.com

A Starting Comment

Remember, it’s your trip to Disneyland we’re talking about. Feel free to ignore any and all recommendations in this or any other guide. Personally, my recommended use for this guide is to help you visualize and plan your trip, feed your dreams of Disneyland, and give you a rough idea of what you want to do. I don’t recommend taking it with you into the park; instead, leave all the guidebooks behind and go have fun.

But then, that’s my idea of how to have fun at Disneyland. I tend to go through Disneyland like there’s no tomorrow, getting there when it opens and leaving when it closes. You may want to do things differently. A pre-planned itinerary may reduce the amount of time you spend waiting in line, but at the expense of spontaneity. The Unofficial Guide to Disneyland by Bob Sehlinger (see Other Resources) gives some itineraries that, I have heard, save a lot of time. I’ve never used them, though, so I can’t tell you how well they work.

Definitions

Here are a few definitions of terms that I may use, but which you may not recognize:

Audio-Animatronics: the art and science of making inanimate objects come to life. This is the technique used on many Disneyland rides to make bears sing, dolls dance, and abominable snowmen scream. It is called Audio-Animatronics because the programs that control the characters are on the tape with the sound, thus ensuring that they never get out of sync.

berm: The earthen wall that surrounds the park. The Disneyland Railroad runs along the top of the berm. There are a few places where passing through the berm is possible. The current list is:

— The main entrance

— The entrance to Toontown

— The middle of Critter Country

— The finale of Splash Mountain

— The Haunted Mansion

— Pirates of the Caribbean

— The monorail

— The entrance to the Indiana Jones ride.

— It’s A Small World

cast member: An employee of the Walt Disney Company, especially one who deals with the public. In this guide, the term “cast members” will refer to Disneyland employees who deal with the guests.

Disney-phile: A person who is fanatical about Walt Disney, his works, and the company named after him. There are many locations in Disneyland that cater specifically to Disneyphiles, such as the Disneyana shop, the Disney Gallery, and the Walt Disney Gallery.

guest: A customer of the Walt Disney Company. In this guide, this will usually refer to visitors to Disneyland.

Magic Kingdom: Disneyland.

mountain range: All the “big” roller-coaster-style rides at Disneyland are mountain-themed in one way or another. Because of this, when Splash Mountain was added, the advertising referred to adding another mountain to Disneyland’s mountain range. The Matterhorn is the oldest, followed by Space Mountain, Big Thunder Mountain, and Splash Mountain.

original lands: Disneyland has grown quite a bit since its opening. This growth has included both adding to existing lands and the creation of entirely new lands. The original lands are Main Street, USA, Adventureland, Frontierland, Fantasyland, and Tomorrowland; the newer lands are New Orleans Square (added in 1966), Critter Country (added in 1972 as Bear Country, then renamed in 1988), and Mickey’s Toontown (added in 1993).

Planning The Trip

Planning the trip in advance is highly recommended. By getting these details handled before you arrive at Disneyland, you can significantly reduce the hassles you have to deal with at the park.

Hotels and Travel Arrangements

I have no firsthand experience with any of the hotels in the area. However, there are quite a few hotels there; at least one is likely to be in your price range. Travel arrangements can be made through your local travel agent, or through the Walt Disney Travel Company at (714) 520-5050. Some of the closest hotels are:

The Disneyland Hotel The Official Hotel of Disneyland. This hotel is connected to the Magic Kingdom by monorail, and, as part of the Disneyland Resort, offers all the attention to detail you expect from Disney. It also has the prices you expect from Disney: expensive. Souvenirs you buy in Disneyland can be delivered directly to your hotel room if you stay here.

Pan Pacific Next door to the Disneyland Hotel; it has a walkway providing access to the Disneyland Hotel’s shops and monorail. Rates are approximately half that of its neighbor.

Most major hotel chains have at least one, if not more, hotels in the area. (According to the map in Sehlinger’s, there are 10 Best Westerns in the area, along with 4 Travelodges, 3 Holiday Inns, a Ramada, a Hyatt, a Hilton, a Mariott, a Howard Johnson’s, and many, many others.) More information on local hotels can be obtained from the Anaheim Area Visitor and Convention Bureau at (714) 999-8999.

Delta Airlines is the Official Airline of Disneyland; however, flights to Los Angeles International (LAX) or the Orange County airport (John Wayne International) can be arranged on virtually any major airline. While flights to LAX may be easier to come by, John Wayne International is much less crowded and much closer to Disneyland. Shuttles to and from the airports are provided by most hotels around Disneyland.

Getting To Disneyland

Once you’re in the Los Angeles area, getting to Disneyland isn’t very difficult. By car, Disneyland is just off Interstate 5. Coming in from the north, you can tell you’re almost there when you can see the Matterhorn off to your right. Take the Harbor Blvd. exit, and go south on Harbor to Disneyland. Look for the large sign on the right, less than a block from the freeway. Coming from the south, ignore the signs telling you to get off on Katella, one exit earlier. Go to the Harbor Blvd. off ramp, take a left onto Harbor, and you’re there!

If you’re staying in a hotel in the area, it probably has a bus or tram to get you to Disneyland. If that doesn’t work, and you don’t want to walk or drive over, Orange County Transit is a relatively good, inexpensive method of traveling. Check the schedules, however, to make sure there is a bus going back when you want to leave; they may not run as late as you’d like to go. Orange County Transit works especially well if your party is going in two groups; the first can ride in on the bus, while the second takes the car (or vice versa). Arrangements are made to meet at a specific time at a specific place, and everybody rides home in the car.

Paying for Things

Disneyland accepts Visa, Mastercard, American Express, and the Disney Credit Card for admission and at shops and sit-down restaurants. Cash is accepted everywhere, as are “Disney Dollars,” Disneyland’s own currency. Disney Dollars can be purchased at Disneyland or at any Disney Store at a one-to-one exchange rate to U. S. dollars. Foreign currency is accepted at most stores, or can be exchanged at the Penny Arcade or the Starcade.

Checks can be cashed at the Penny Arcade on Main Street or in the Starcade in Tomorrowland. Most stores will also accept checks for the amount of purchase. ATM machines are available outside the main entrance (to the left, as you face the gate), at the Bank of Main Street, in the Frontierland stockade, in the Premiere Shop in Tomorrowland, at the Penny Arcade, in the Starcade, and in the Disneyland Hotel.

Disney Dollars can be a good idea, or they can be hazardous to your financial health, depending on how you use them. If you know how much money you can spend at Disneyland, you can buy that much in Disney Dollars and spend only those. However, since they don’t look like “real” money, they’re usually easier to spend.

A Magic Kingdom Club card is a very useful thing to reduce the amount of money you spend at Disneyland. A card will drop your entry prices by 5 to 10 percent, and may give you a discount on food inside the park. Your employer may offer memberships; many large companies do. If you can’t get a membership this way, they can be purchased for $59 for two years. Application forms are available at most Disney Stores.

In addition, AAA membership will get you a discount on the price of admission.

Getting Ready To Go

Start the day early, with a good breakfast. A good breakfast is important so that you can wait until after the noon rush to eat lunch; peak times for lunch are from about 11 AM to 1 PM. You may want to get breakfast outside the park; food can be expensive inside Disneyland.

Appropriate attire is required for park entry. Bikini tops, skimpy clothing, or clothing bearing offensive words or pictures are all strongly discouraged by Disneyland, and you may be refused entrance until you’re dressed more appropriately.

Wear comfortable clothes, especially shoes. You’ll be doing a lot of walking and spending a lot of time on your feet. Disneyland is not a good place to break in your new sneakers, or to walk around in high- heeled shoes. Even if you’re going in the hottest part of the summer, you may want a sweatshirt or jacket in the evenings. Bring them along and either leave them in the car, store them in a locker, or bring a backpack. Most rides have places where a backpack can be accommodated, usually on the floor in front of you. Plus, it gives you a place to put all the stuff you buy at the shops.

If you want you can bring a light, non-messy snack, like a bagel, into the park in this backpack. Most cast members won’t have a problem with this as long as you don’t point it out to them. Technically, however, food should not be brought into the park. If you want to bring your own food, there are lockers outside the park which are designed to hold coolers; the lockers are not refrigerated, however. There is a picnic area just past the lockers.

Sunblock, film, batteries, and other miscellaneous items are much cheaper to bring in than to buy in the park. Plan to bring them in with you to save yourself time, money, and hassle.

Annual Passes

Disneyland currently has two kinds of annual passes available, the basic annual pass and the premium annual pass. The basic annual pass costs $99 ($89 with a Magic Kingdom Club card) plus an optional $25 parking fee. This includes a free run of the park most days, with exceptions being mostly Saturdays from February to June, and certain holidays. If you go more than three or four days in a year, you will certainly save money with a pass.

The premium pass costs $199 ($189 with a Magic Kingdom Club card) and gives unlimited entrance every day of the year, free preferred parking, a 10% discount on merchandise and food (except walkup windows) inside the park, and a free one-year subscription to Disney Magazine. If you buy a lot of merchandise inside the park, this card can easily pay for itself.

Kids

I don’t know very much about going to Disneyland with kids; for those travelling with children, any of the published guidebooks will have details, especially Disneyland with Kids by Wiley. The comments I’ve received from the guidebooks and from people who have gone with children generally indicate the following:

How you plan your trip will depend largely on the ages of your children. For example, a three-year old might be best with a short day in Fantasyland and Toontown, while a twelve-year old might want to stay all day and hit every ride.

Check the height and age requirements for the rides ahead of time. There are signs at the main entrance of Disneyland which list the attractions with restrictions, and what the restrictions are. At the time of writing, children must meet these restrictions to ride the following rides:

— Autopia (1 year to ride with an adult, 52 inches to ride alone)

— Star Tours (3 years with an adult, 7 years alone)

— Matterhorn Bobsleds (3 years with adult, 7 years alone)

— Gadget’s Go-Coaster (3 years old)

— Big Thunder Mountain (40 inches)

— Space Mountain (40 inches)

— Splash Mountain (3 years and 40 inches)

— Indiana Jones (3 years and 48 inches)

Make sure your children realize which rides they won’t be able to go on before you get to the park.

Realize that children don’t usually have the endurance that adults do. Prepare to take time out for naps, rest, bathroom breaks, and food.

Stay close to the park. If you don’t want to stay at the Disneyland Hotel, stay at one of the other hotels close to the park. This will let you go back to the hotel room for a break at midday if necessary.

Keep an eye on young children. Disneyland can be a very exciting place for them, and they may run off to look at something while you’re looking in another direction.

Lost Children: If you and your child become separated, Lost Children registers are kept at City Hall and Lost Children, next to First Aid at the Central Plaza end of Main Street. Any cast member can give directions to either of these places. Disneyland is very good at reuniting parents and children quickly.

Disneyland – Park Basics

You’ve planned everything you’re going to plan, brought everything you’re going to bring, and now you’re getting ready to go into the Magic Kingdom itself. Here’s some basic park information.

Outside The Main Gate

Outside the main gate are various services for guests, such as:

a Guest Relations center

restrooms and public phones

a souvenir shop and newsstand

an ATM which accepts virtually any ATM card

a picnic area

a kennel to keep your pets in while you’re in Disneyland

storage lockers

Getting In

There are two ways into the Magic Kingdom. The most commonly used of these is through the front gates. This is the easiest way in for people coming by car, as the front gate faces the parking lot. Also, most of the buses from other hotels drop off and pick up in the parking lot.

The other way into the park is on the monorail. You get on the monorail at the Disneyland Hotel, across the street from Disneyland, and get dropped off in Tomorrowland. It is most convenient for Disneyland Hotel guests and people walking from hotels to the west.

So which way should you go in? Go in by whichever way is most convenient for you; each gives its own perspective on the park. The main entrance starts you in Main Street, which leads you into the central area of the park, while the monorail gives descriptions of some of the areas it goes over and starts you in Tomorrowland, right near some of the biggest rides in the park.

One factor that can decide how you enter is whether you buy tickets ahead of time. If you plan to buy your tickets at the gate, go to the main entrance. While there are many ticket booths at the main gate, there is only one at the monorail station. The long line for buying tickets at the monorail station can be bypassed if you already have your tickets.

Meeting Places

If you’re going as part of a group that doesn’t plan to stay together all day, arrange times and places for at least your first meeting as soon as you get there. In addition, arranging a meeting place in case you get separated by accident is a good idea. Some good places to meet include:

The Central Plaza, except before and during parades. There are benches and a lot of room here, and, if you pick out a specific statue to meet near, it’s easy to find everyone else. For an hour or two before a parade, however, it’s a mess.

The entrance to Tomorrowland. There’s a small “island” in the middle that makes a good waiting spot.

The waiting area for CircleVision, especially if it’s a little chilly or very hot out. Use this one only during non-peak times, and don’t plan to use it right before closing, because it closes a bit before the rest of the park.

The Snow White Grotto at the side of the castle. (Never crowded, and hard to confuse once you know where it is. It’s on the side of the castle closest to the Matterhorn, at the end of the moat.) Again, it’s a bit difficult to get to during a parade, but doesn’t have the problems before the parades that the Plaza does.

The lower level of the Hungry Bear restaurant in Critter Country. While this isn’t as central as the other waiting areas listed, it has a roof and convenient restrooms, and you can just go upstairs to get food. This is a particularly good meeting place for a late lunch, since the Hungry Bear’s lunch rush ends early compared to the rest of the park.

In front of the castle is not a good meeting place, no matter how impressive it looks. Since it looks so obvious, everyone who’s never been to Disneyland before tries to meet in front of the castle. Add in the traffic between Fantasyland and the central plaza, and it’s impossible to find anyone.

Be sure your entire party knows where the meeting spot is, preferably by going there before you split up. This ensures everyone is thinking of the same place. It can be very easy to get confused about exactly which tree, post, or store you were supposed to meet at if you haven’t walked over to it and said “We meet right here.” This also prevents getting to your location only to find that it’s closed for refurbishing.

Lines

One of the most common things found at Disneyland is the line, or queue. There are lines for virtually everything: rides, food, cash registers in the shops, getting your picture taken with Mickey Mouse, and more. However, a line at Disneyland is quite unlike a line anywhere else. Readers who haven’t been to Disneyland may feel shocked when I refer to a twenty-minute line as short; however, a twenty-minute line at Disneyland feels shorter than a five-minute line at the grocery store.

There are a few reasons for this: first, people are on vacation, so they’re more relaxed; second, Disney gives you interesting things to look at on each line; third, the lines move very quickly; and fourth, Disney does everything it can to keep you from seeing the whole line at one time. There is always another bend for you to watch yourself approach; always something to catch your eye; always some separator for you to pass.

Even a long line is not a terrible experience. The most noticeable thing about my last 60-minute wait for Splash Mountain was that my feet were starting to hurt, and that was due more to the fact that I had been going through the park non-stop for twelve hours.

A few tips on avoiding the worst lines:

Hit the attractions with big lines either very early or very late. Which works best depends on the ride; for example, Star Tours is best done late, as are most Fantasyland and Toontown attractions. Space Mountain, on the other hand, is usually easiest to get on very early.

At midday, when lines are at their worst, visit attractions that are either less popular or have a large capacity. Examples include the Country Bear Playhouse, The Walt Disney Story, The Enchanted Tiki Room, and the Haunted Mansion.

After you’ve seen the parades and shows, use the times when other people are watching them to see attractions.

As a general rule, the point where a line enters the building is approximately 10 to 20 minutes from the ride.

If you can, go on a day when there will be fewer people. The busiest days are Saturday, then Friday and Sunday. Monday through Wednesday are the slowest days. The disadvantage to this is that during the off-season, the park has shorter hours, especially on weekdays, and nighttime shows may not be run.

On busy days, lines die down after the first Fantasmic! performance, because many families leave. After you see Fantasmic!, visit the high-profile rides, where the lines should be very short to nonexistent.

Disneyland always overestimates the time to get to the ride; the listed time is, usually, the absolute maximum you can expect to wait, barring technical difficulties.

Restaurant Reservations

Reservations are highly recommended, although not required, for easting at Aladdin’s Oasis or the Blue Bayou. Reservations are required (except during very slow seasons) for the Disney Gallery’s dessert buffet and private seating for Fantasmic!, as well as for the “Billy Hill and the Hillbillies” show at the Golden Horseshoe.

Reservations must be made at each respective location, preferably as early in the day as possible. There is a phone available in City Hall for making reservations at the restaurants in the Disneyland Hotel.

Performances

Every day, Disneyland puts on a show for its guests. On weekends and during the busy season, there’s at least two parade performances (currently The Lion King Celebration), and when the park is open late, there’s usually at least two showing of Fantasmic! During the summer, the Main Street Electrical Parade and Fantasy in the Sky Fireworks are also performed. For more details on the current shows, call Disneyland Guest Relations at (714) 999-4565 or check Walt Hoffman’s Disneyland Frequently Asked Questions list or Al Lutz’s Disneyland Information Guide.

Schedules of a given day’s performances are listed in the “Inside Disneyland Today” pamphlet, available at the Main Entrance, City Hall, and many of the major shops.

Parade This is a parade through the Magic Kingdom, starting at either Small World or the Town Square, and proceeding to the other one. The current parade is The Lion King Celebration. Curbside seats may give youngsters an opportunity to play an instrument during the show; getting to Town Square early is probably the best way to get a curbside seat.

Main Street Electrical Parade During the summer, the nights are filled with this entertaining parade of brightly lit floats and performers, traveling the normal parade route. It’s fun to watch, and has a bouncy soundtrack. Performed twice nightly during the summer, this is a long-standing, popular Disneyland tradition.

Fantasy in the Sky Fireworks This is perhaps the most famous show Disneyland puts on, as every Friday and Saturday night (every night during the summer) fireworks light up the sky above Sleeping Beauty Castle. It is performed right after the first Electrical Parade during the summer, and marks the end of the day for many families. After the fireworks, the park gets noticeably less crowded. Try to get a view of the Castle or Matterhorn to see Tinker Bell “fly” past to “light” the fireworks!

Light Fantastic The latest addition to Disneyland’s nighttime shows is the tentatively-titled “Light Fantastic” featuring 25 “searchlights” placed around the park. Each light projects a beam of colored light into the sky during the fireworks, backlighting the castle at the end of the show. It’s worth being on Main Street to see this! The “Light Fantastic” display is also done throughout the rest of the evening, for about five minutes at a time.

Fantasmic! Don’t miss this spectacular water, pyrotechnics, light, and stage show, starring Mickey Mouse on a journey through his imagination. With appearances from many favorite Disney characters, this show has become one of the most popular shows in the park. It’s usually done any night the park is open late, and is very highly recommended. There are some strobe-type lighting effects in the show, so guests with physical conditions which may be adversely affected should use caution.

Fantasyland Theater The Fantasyland Theater, (formerly named Videopolis), near Toontown, frequently has stage shows. The current show is ‘Spirit of Pocohontas.’ It plays a number of times a day, from afternoon through evening. Be sure to arrive early to line up for a good seat.

Disneyland’s 40th Birthday Celebration Throughout the 40th birthday year of 1995, there is a brief birthday celebration in front of the castle at noon. Disney characters will be there, and will sign autographs afterward.

To get a good spot for a parade, get there early. A good place to see most of the parades from is the central plaza; unfortunately, many other people realize this too. Another good spot, especially for the Main Street Electrical Parade, is in front of “It’s A Small World.” The parade route runs much closer to the viewers here than in other places. Staking a spot out 60 minutes or more before the parade starts is not uncommon, especially along Main Street and in the central plaza. Town Square, however, is a great spot to find good spots available.

To get a good spot for Fantasmic!, I recommend waiting until the second show if possible. Many people with younger children will watch the first show, then leave. To get a good spot, head over to Critter Country before the first show starts, and maybe hit Splash Mountain and/or the Country Bear Playhouse. Stay back in Critter Country until about twenty minutes after the show starts (it lasts about twenty-three minutes), then head in as the first show’s audience gets herded out through Frontierland and Adventureland. If you get the timing right, you should get seats in the front-most section. Be warned, you may get misted from some of the water effects if you’re right in the front. You should have about a forty-five minute wait for the next show; a good coat is recommended, since it can get a little chilly. One of your party can run over to La Petite Patisserie in New Orleans Square for hot chocolate and coffee, or to the Royal Street Veranda (near the entrance to “Pirates of the Caribbean”) for coffee, Mardi Gras punch, and fritters, or to the Mint Julep bar in the French Market seating area. (Note that the mint juleps are non-alcoholic.)

On the other hand, if you want to do your rushing around in the morning and don’t mind spending some money, the Disney Gallery has “reserved” balcony seating for Fantasmic! Make your reservations very early, at the Gallery, then, for $30, get an all-you-can-eat dessert buffet and prime seating to watch the show.

If you’re not going to be watching a particular parade or show, use it to your best advantage. Usually, a big parade or show will draw people away from the rides, allowing you to wait in a much shorter line. Plan for this, especially for parades; crossing the parade route can be difficult. If you get stuck on the “wrong side” of the parade route, your best bet for crossing is often the Railroad, although Disneyland makes every effort to allow guests to cross whenever it is safe to do so.

Meeting Characters

The best places to meet characters are Town Square, in front of the Castle, and, especially, Toontown. If there’s a particular character you want to meet, ask at City Hall; they can tell you where that character can be found. On hot days, look for the characters in shaded areas, sometimes even under large picnic-sized umbrellas.

A new method of handling characters (with formal lines) is being tested in New Orleans Square (in front of the Haunted Mansion) and in Toontown. By controlling access, Disneyland helps keep the characters, and the children around them, from getting mobbed, while giving Mom and Dad even better opportunities for taking pictures and videos.

Introducing children to the Disney characters can cause a few problems if a few factors aren’t taken into account by the parents. Things parents need to consider are:

The characters are a lot larger than you’d think. Many small children looking forward to meeting Mickey have been terrified on realizing that Mickey is close to six feet tall. In general, take the children to meet the characters toward the end of your visit, when they’ve had a chance to get used to their size.

Most of the characters don’t talk. However, they’re usually extremely good at getting their point across without words.

The characters may not be able to see small children that are directly in front of them. Due to the designs of the outfits, especially those for Donald and Daisy Duck, their field of vision may be limited. Lift the child to the character’s eye level, and they’ll make sure the child gets attention.

Be polite to the characters. They are an especially magic part of the Magic Kingdom, and don’t deserve the assorted abuse they’ve received over the years.

Cast Members

Cast Members in Disneyland are, in general, friendly, helpful, and willing to do anything reasonable to ensure the happiness and safety of the guests. Most cast members work at Disneyland because they like Disney and the Disney spirit, and it shows. Feel free to ask any cast member for assistance with any problem you may be having at the park, but please don’t try to make their lives difficult. They are a very large part of what makes Disneyland the wonderful place that it is.

In particular, cast members are almost always willing to take pictures of your group, so that everybody can be included in the shot. However, if you are planning to do this, take a “point-and-shoot” camera; while they’re willing to take the picture, they need to be able to use the camera to do it.

Any compliments or complaints about cast members can be formally entered at City Hall, and action will be taken on each.

Dealing with the Real World

Despite the best efforts of Walt and those who have followed him, the real world does intrude into Disneyland somewhat. Some things that you may need to deal with include:

Rain. It does rain on Disneyland occasionally. Rainy days are actually some of the best days to go because the lines are much, much shorter than normal. Rain ponchos and umbrellas are sold at the Emporium on Main Street, among other places. You have to ask for them, usually, because they’re kept behind the counter except on very rainy days.

Car problems. If you get back to your car at the end of the day just to find that your battery is dead, find one of the parking lot security people. They’ll put you in contact with someone who can help. If your problem is that you can’t find your car, security can send someone out on a scooter to help locate your car. Don’t hesitate to ask–that’s what they’re there for!

Illness. Aspirin and other common medicines can be found at the Emporium on Main Street. Ask at the counter for them, because they aren’t put out on display. In addition, the First Aid center on Main Street has a registered nurse on duty at all times the park is open. If you feel acutely ill, ask any cast member to get the nurse, and help will arrive quickly.

Walt’s Touch

Part of Walt Disney’s beliefs about Disneyland could perhaps be summed up as there being no such thing as an insignificant detail. Every corner of Disneyland shows this belief, from the lights and landscaping to the trashcans. Some of the details to look for include:

The names on the second-story windows on Main Street are people who contributed to the design, implementation, and running of Disneyland. One of the most recently added is for Jack Lundquist, the recently retired President of Disneyland.

Above the Fire House on Main Street is Walt’s apartment, which he used when he stayed overnight at Disneyland. A small lamp in the window is kept constantly lit, in his memory.

Main Street and Sleeping Beauty’s Castle were both built with the use of “forced perspective.” They were built so that the ground levels were at normal size, but higher areas were progressively smaller. This makes the upper areas look even higher when seen from below.

Gum is not, and never has been, sold inside Disneyland. Any gum found on the ground is scraped up immediately. Walt didn’t want his guests to have to pick gum off their shoes every ten minutes, as was (and is) common at other amusement parks. Alcohol is also not available in any public area of the park.

The horses on the King Arthur Carousel are all individually hand- carved and painted. They are continually being repainted, and all the brass is polished every night.

Walt Disney was very insistent that the designers take the perspective of young children into account when designing the park. Because of this, Walt would frequently stoop down while looking at a partially constructed building to take into account how smaller people would see things.

Disneyland’s address is 1313 Harbor Boulevard. This address was picked by Walt, because at the time Disneyland was built, all that was in the area was orange groves. The explanations for the address are either that ’M’ is the thirteenth letter in the alphabet, making the address MM for Mickey Mouse, or that the thirteens, being unlucky, were Walt’s way of thumbing his nose at those who thought Disneyland would fail.

Most gold trim in the park is done with real gold leaf rather than gold paint. This is partially because the gold leaf looks spectacular, and partially because the gold leaf may actually be cheaper because it doesn’t need to be replaced as often.

When New Orleans Square was added, Walt wanted to put in a private club at which he and other dues-paying members could entertain. Sadly, he was not able to see it completed, but the club was finished and is now a members-only club that occupies much of the second floor of New Orleans Square. The club, called Club 33 for its address, 33 Rue Royale, is the only place in the park where alcohol is served. The membership fee is (at the last time I heard) $5000 to start plus annual dues of $1800. There is a waiting list to become a member; the latest information that I’ve seen said that the waiting list was about two years long.

Walt made sure the grass areas weren’t fenced in until he saw which areas people walked on. Areas that were used frequently as short-cuts had sidewalks added.

The petrified tree in Frontierland was an anniversary present that Walt gave to his wife, Lillian. She donated the tree to Disneyland, where it was given its spot overlooking the Rivers of America.

The Jungle Cruise got a major overhaul based on what Walt overheard one guest say. The remark in question was: “We don’t need to go on this ride, we’ve already seen it.”

Special Assistance

Disneyland has tried to make as many attractions as possible wheelchair-accessible, and gives out lists of which attractions have wheelchair access. Wheelchairs can be brought in or rented inside the park, right by the entrance with the stroller rentals. Small, one-person electric carts can also be rented.

A Disabled Guest Guide is available from Guest Relations or City Hall with specific information on accommodations for disabled guests.

City Hall also has guide brochures in foreign languages, as well as printed text of some attraction narrations, for the benefit of guests with hearing disabilities.

Disneyland – The Lands

To help give each area of the park its own theme, Walt divided the park into ‘lands,’ each of which has its own characteristic appearance. The lands added since Disneyland opened have kept this up, and each land’s theme can be seen in the style of building, in the cast members’ costumes, and even in the appearance of the trashcans.

Main Street, USA

Once you enter the main gate, there is a small plaza before you pass through the berm into Disneyland. In the plaza, there is a portrait of Mickey made of the flowers on the side of the berm. Above the tunnels through the berm are plaques, which read:

Here you leave today and enter the world of yesterday, tomorrow, and fantasy.

After going through the tunnel, you emerge into the Town Square on Main Street, USA. Main Street is one of the original lands, and is modeled after the town Walt grew up in. It is also the shopping mecca of Disneyland. Almost anything you want with a Disney character on it is available somewhere on Main Street.

Some of the stores available on Main Street include: the Mad Hatter (just the place to get embroidered mouse ears!), a candy store, a storybook store, a watch and clock shop, a crystal store, a magic shop, a Disneyana shop, a photography studio, clothing stores, the Showcase store (featuring items from the latest animated features and rides), various snack stores, and Emporium, which has more than a little bit of everything.

There’s more than just shopping on Main Street, though. It is also the administrative center of the park. The Disneyland City Hall and Guest Relations are both on Main Street, as are the lockers, Lost and Found, First Aid, the Baby Care Center, and the Lost Children area. The Baby Care Center, in particular, is a service many people are unaware of. Tended by nurses, the Center provides a place to nurse babies, warm formula, and change diapers. Rocking chairs are provided for mothers and mothers-to-be.

Annual passes are processed here, in the Bank of Main Street, next to the Walt Disney Story. Guided tours start in a small garden off the town square, and must be purchased with your entrance ticket in the morning.

Main Street also has many restaurants and places to buy snacks. The Carnation restaurant provides table service and a varied menu. The Cone Shop, near the lockers, has ice cream creations and drinks. Refreshment Corner (commonly called Coke Corner) has hot dogs, pretzels, and drinks. The Blue Ribbon Bakery has cookies and frozen yogurt, and the Plaza Inn and Plaza Pavilion buffeterias offer more substantial meals. Other, smaller shops sell snacks from fruit to cookies and hot chocolate.

This are is somewhat busy in the early afternoon, and extremely busy around closing time. Because so many people wait until they’re leaving to do their shopping, Main Street closes later than the rest of the park. Waiting until the last minute to shop is a bad idea, however, as the stores are mobbed at closing time. Instead, do your shopping earlier in the day, and either carry it with you, store it in a locker, or, have it sent to the Package Pickup center to be picked up on the way out. This is a service provided by most shops in the park; you’ll need to allow a few hours for your purchases to make it to the pickup center, but it’s a lot easier than carrying everything with you.

Important attractions and locations on Main Street include:

Main Street Vehicles Horse-drawn streetcars, horseless carriages, a fire wagon, and the Omnibus (a double decker bus) are all ready to take you from the town square to the central plaza or back. There are rarely lines, but waiting for a ride is rarely faster than walking.

The Walt Disney Story featuring Great Moments with Mr. Lincoln This is a pleasant walkthrough about Walt Disney and the building of Disneyland, followed by an Audio-Animatronic Abraham Lincoln delivering a speech. The lines are very short, and the show is relaxing. Featured items on display include Walt Disney’s working office and formal office, both relocated exactly “as was” (including carpeting) from the studios in Burbank. Younger children may get bored, so this is probably not a very good attraction for them, but for Disney-philes, this is not to be missed.

Main Street Cinema Six short cartoons run simultaneously, including the first “talkie” cartoon, a Mickey Mouse short entitled “Steamboat Willie”. There are no lines, and the theater is air-conditioned.

Disneyland Railroad This ride features a trip around the edge of Disneyland. The railroad stops in three other places around the park: New Orleans Square, Toontown, and Tomorrowland. The locomotives are steampowered, and run along the top of the berm. Most of the seats in the cars face inwards, towards Disneyland, while a few face forward. In many ways, the trains were the first part of Disney’s plan, which began with the idea of putting a train around the Disney studio in Burbank. On the ride around the park, you get to see parts of some of the other attractions, such as Splash Mountain, as well as some views only visible from the railroad, such as the Grand Canyon diorama and the Primeval World dinosaurs. One of the engines, the Fred Gurley, recently had its 100th birthday!

Package Pickup Center Guests may have their purchases held for them at the package pickup center, located in the Town Square next to the Penny Showcase. When you make a purchase, ask the cashier for details about having your packages held. If you’re staying in the Disneyland Hotel, your packages can instead be delivered directly to your room.

Central Plaza: the Hub of Disneyland

The Central Plaza (or, as it’s commonly called, the Hub) is technically considered part of Main Street, but is in many ways separate. A recent addition to the plaza is a beautiful, life-size bronze statue of Walt and Mickey standing hand-in-hand, smiling and looking down Main Street. The statue is titled, appropriately enough, “Partners,” and has a quote on the base:

I think most of all what I want Disneyland to be is a happy place… where parents and children can have fun, together. – Walt Disney

Around the statue is a small flowerbed, along with benches and some smaller statues of other Disney characters. This area makes a decent meeting area, as long as you aren’t meeting at parade time.

Just past the central plaza is the front of Sleeping Beauty castle, perhaps the most recognizable symbol of Disneyland. Passing through the castle leads you to Fantasyland; each of the other original lands is also accessible from the central plaza, which acts as the “hub” of Disneyland. Look down at the water as you pass over the bridges, and you may see fish, swans, and ducks swimming around underneath you!

A new addition to the central plaza is a chalkboard listing the current wait times for various popular rides, as well as show and performance times. Check it out; it should give a good idea of where the crowds are.

Adventureland

Adventureland is one of the original lands, and is also one of the smaller lands as far as attractions are concerned. However, the attractions in this land make up for its lack of size. The buildings are in a combination of African, Middle Eastern, and Polynesian styles, making the whole setting reminiscent of an exotic, faraway place. Included in this land are many shops with exotic-type goods, as well as restaurants like Aladdin’s Oasis (open for lunch and dinner during peak seasons; make reservations here early).

Adventureland’s attractions include:

The Enchanted Tiki Room This is an Audio-Animatronic show with singing birds and catchy tunes. Lines are rarely bad, and the theater is air-conditioned. Catch it when lines are bad elsewhere during the day. Older children (that is, retirement age and up) seem especially fond of this attraction.

Jungle Cruise The Jungle Cruise is a voyage through the jungle with many humorous views and wisecracking (human) tour guides. Residents include elephants, hippos, snakes, and headhunters. This is one of the few attractions still in existence that was around on opening day. It’s great fun, and is famous for the corny jokes told by the guides. Lines aren’t usually very bad, now that the Indiana Jones attraction has opened next door. Walt originally wanted to put live animals in here, but decided to put in the Audio-Animatronic ones after being informed that the animals would be sleeping most of the time that guests would be around.

Swiss Family Treehouse This large tree, with stairs leading up to the home of the Robinson family, towers over Adventureland. This is a walk-through attraction, with views of many wonderful devices the family created to help them survive. It’s entertaining, and rarely very crowded. The tree is an exceedingly rare Disneydendron, with over 300,000 artificial leaves. Smaller children tend to want to take their time going through this and look at everything, but larger children (that is, past high school) tend to get hurried along by the groups behind them. Attentive guests will notice that the tree home was established on July 17, 1805, exactly 150 years before Disneyland opened on July 17, 1955.

Indiana Jones Adventure – Temple of the Forbidden Eye Disneyland’s newest attraction, this expedition through the Temple of Mara combines the track of a “tracked car” ride like Mr. Toad, the hydraulic motion-control technologies of Star Tours, and clever computer programming that changes the ride every time through. Add in a state-of-the-art onboard sound system, and you get a thrill ride like no other. This attraction is very highly recommended. However, it has the highest height requirement in the park, 48 inches, so smaller children won’t be able to ride. The sign says that the ride is not recommended for people with back, neck, or heart problems, or for pregnant women; they mean it. This is a seriously intense ride. Lines for this attraction, as the newest in the park, are relatively long, with 60 to 90 minutes being the usual length. Lines have been known to reach up to three hours long at peak times on busy days. The lines for this attraction start on Main Street before the rest of the park actually opens. Also, since this ride is new, technical difficulties are more common than other rides. Get in line early, and plan to wait a bit, but it’s worth it. It’s also a good idea to check the line late in the evening, after most of the crowds have gone home. The queue area is designed to keep guests entertained while they wait, so the line doesn’t seem as bad as it could be.

New Orleans Square

New Orleans Square was the first land added after the park opened, and the last land Walt worked on personally. It is one of the busiest lands in the park, mostly because its attractions are among the most popular in the Magic Kingdom. There are a few shops, and some restaurants, including the Blue Bayou, which overlooks the beginning of the Pirates of the Caribbean ride.

This land is also the home of Fantasmic!, the spectacular water/light/action show that takes place on the Rivers of America at night. For more details, see the Performances section.

Attractions in New Orleans Square include:

Pirates of the Caribbean This is quite possibly the most popular ride in all of Disneyland, and deservedly so. The lines can get a little long during the peak hours, but are usually quite reasonable because so many people can ride at once. This is a ride through a pirate raid on a Caribbean island, and is an excellent examples of Audio-Animatronics at work. It begins with a view of the hidden pirate treasure, and carries you through battles and bride auctions to the destruction of the town. The theme song, “Yo Ho,” is one of the best-known songs in all of Disneyland. This is the last attraction Walt Disney personally supervised before his death.

Disney Gallery This art gallery is located right above the entrance to Pirates of the Caribbean. The current displays include “The Art of Disneyland,” “The Art of Disneyland Paris,” and much of the artwork created during the design of the Indiana Jones attraction. Featured works include Herbert Ryman’s original watercolor of Sleeping Beauty Castle, Peter Ellenshaw’s aerial view of the “proposed” Disneyland, and the scale model of the castle used in planning Disneyland. While the historical pieces are not for sale, many of the other pieces on display are. In addition, the Gallery also maintains an extensive collection of animation cels, available for purchase. A must-see for Disney-philes.

Haunted Mansion Meet up with 999 ghosts; there’s room for a thousand, in case you’d care to join them. This attraction, while very popular, usually has relatively short lines because many people can ride at once. The ride is not particularly scary; it is, however, very entertaining. You begin in a small room, which leads you deep into the mansion’s hidden depths. Make sure you read the tombstones in the “pet cemetery” that the line winds around. Some young children may find this ride too dark and too scary.

Disneyland Railroad The railroad makes a stop in New Orleans Square on its passage around the park. The ride from New Orleans Square to Toontown includes a look at the Rivers of America and the finale of Splash Mountain.

Frontierland

Frontierland is Disney’s homage to the “Wild West” of America’s past. The buildings look as if they came out of a cow town in Texas, the shops sell Western apparel, and there is a shooting gallery (coin-operated, and shooting infrared beams instead of bullets).

Attractions in Frontierland include:

Big Thunder Mountain This ride is the tamest part of Disneyland’s mountain range of roller coaster-style rides. There’s nothing more severe than drops and sharp turns, and you may get a few drops of water on you, but nothing more. Audio- Animatronics are used well on this ride, bringing to life such animals as goats, vultures, prairie dogs, and lizards. There are two ways to enjoy this ride: go during the day for full effect of the local fauna, or after nightfall for a more intense ride in the dark. Lines are rarely very bad, with twenty to thirty minutes being typical. There are boarding restrictions: riders must be at least 40 inches tall, and expectant mothers and those with back, neck, or heart problems are advised not to take this ride.

Rivers of America This is a wide waterway around Tom Sawyer Island. These waters are also used for the Fantasmic! show at night.

Steamship Mark Twain / Sailing Ship Columbia These two boats give tours around the Rivers of America, showing views of frontier life along the river, including a Native American village. The Mark Twain is a five-eighths scale reproduction of a Mississippi riverboat. It was built in two pieces, with the deck built in the Disney studios and the hull being built at a local shipyard. Despite being built miles apart, the two pieces fit together perfectly the first time. The Columbia is a reproduction of the first American sailing ship to go around the world. Walt himself put a silver dollar under each mast, as was customary when building sailing ships. The Columbia was the first three-masted windjammer built in the US in over one hundred years.

Tom Sawyer Island This is an island in the middle of the Rivers of America, where young children can run around and play to their hearts’ content. It’s a good place to let kids burn off some energy after standing in line for a while. The rafts to get over are just across from the Haunted Mansion entrance gate. The island closes at dusk.

Big Thunder Ranch This is Disneyland’s own petting zoo, and is quite possibly the cleanest ranch around. Also out here is Big Thunder Barbecue, a western chow-hall type restaurant serving barbecued chicken and ribs. It’s one of the better deals for the money when eating in the park.

Critter Country

One of the newer lands, Critter Country was originally named Bear Country. The theme here is rustic and natural, and the buildings look as if they’d fit in quite well halfway up a mountain in the middle of the Rockies or the Appalachians. There are only a few shops, and only one major restaurant here: the Hungry Bear, home of hamburgers, hot dogs, salads, and sandwiches. For fans of Winnie the Pooh and friends, the Briar Patch store is devoted exclusively to Pooh items.

Country Bear Playhouse This is a show full of singing bears, corny jokes, and a talking stuffed moose, put on entirely with Audio-Animatronics. It’s enjoyable, the lines are rarely long, the seating capacity is high, and the theater is climate-controlled, all good qualities when you need a break. Go when the lines are at their worst everywhere else. At Christmas time, the bears put on a special holiday show.

Splash Mountain This is the newest addition to Disneyland’s mountain range, and one of the most exciting rides in the park. It’s a log flume ride through part of Song of the South, starring Br’er Bear, Br’er Fox, and Br’er Rabbit and ending in a five-story drop into the “Briar Patch.” When you go over the edge, remember to smile as your picture gets taken by the automatic camera. A warning at the entrance says “You might get wet,” and they aren’t kidding! Hit this ride very early or very late, because the lines are some of the worst in the park; sixty minutes is common, and ninety is not uncommon. However, it makes up for the long lines by being one of the longer rides, easily over ten minutes long. There are height and age requirements for this ride; children must be at least 3 years old and 40 inches tall to go on this ride. People with back, neck, or heart problems and expectant mothers are advised not to ride.

Explorer Canoes The canoes go through the same route as the Steamboat Mark Twain and the Sailing Ship Columbia. Propulsion is provided by the guests, who get to paddle the canoes the whole way.

Fantasyland

One of the original lands, Fantasyland is dominated by the spires of Sleeping Beauty Castle and the towering Matterhorn. Most of the rides in Fantasyland are aimed at younger children, but they’re all fun for all ages. This land gets relatively crowded at midday, since that’s when the most children are around. Because of this, most rides in here should be hit early or late to avoid the rush. There is only one major restaurant, the Village Haus, which serves burgers, pizzas, and salad. The shops include a Christmas store, open year round in the Castle, Tinker Bell’s Toy Shop, with lots of general Disney merchandise, a Heraldry store, where you can find your family crest, and a Villains store, dedicated to the evildoers of the Disney movies.

Sleeping Beauty Castle This castle, possibly the single most recognizable symbol of Disneyland other than Mickey, contains a Christmas Shop, a heraldry store, and a walkthrough telling the story of Sleeping Beauty. The drawbridge is real, and was lowered on opening day.

Snow White Grotto This small grotto, over on the side of Sleeping Beauty Castle, has a wishing well and statues of Snow White, the seven dwarfs, and various woodland creatures. This is a beautiful area to relax for a few minutes, and is never crowded.

Snow White’s Scary Adventures Ride through the story of Snow White, with the Wicked Queen coming for you around every corner. It’s not very scary, but I’ve been told it makes a good ’testing ground’ for the Haunted Mansion for young children. If they can handle this, they can probably handle the Mansion. Look up at the window above the entrance; sometimes, the Wicked Queen opens the curtains to look out upon Fantasyland.

Pinocchio’s Daring Journey Take a trip from Geppetto’s workshop through Pleasure Island and beyond. This is a short, fun ride with very short lines.

Peter Pan’s Flight This ride puts you in a pirate ship, which flies through the Darlings’ bedroom and off into Neverland. When you look down at London, notice Big Ben and the moving cars. It’s one of the best Fantasyland rides, and the lines show it. They are some of the longest in Fantasyland, but still not too long compared to the rest of the park.

Storybook Land Canal Boats The boats carry you through Storybook Land, where you can see the miniature homes of many of Disney’s favorite fairy-tale characters. It’s enchanting, and the lines aren’t bad. Older children (twenties and up) will love the wonderful miniaturization work done throughout Storybook Land.

Casey Jr. Circus Train This is a ride in a circus train, where you can sit in the lion cage or even the caboose. The train track runs parallel to the Canal Boats, and gives a different view on Storybook Land.

Alice in Wonderland Caterpillars carry you through Wonderland, where you meet the most unusual creatures on either side of a looking-glass. This ride is short and entertaining.

King Arthur Carousel This is a very large and detailed merry-go-round, with seventy-two carefully hand-painted horses. Lines are usually not very long. In front of the carousel (towards the castle) is a sword in a stone; at various points during the day, Merlin will give children a chance to pull the sword out (part-way) and become King (or Queen) for the day.

Mr. Toad’s Wild Ride Get in a car named after one of the characters from “The Wind in the Willows,” and ride off on the road to Nowhere in Particular! This ride isn’t especially wild compared to the nearby Matterhorn, but is very entertaining and shows a great attention to detail. Lines are long for Fantasyland, with twenty-minute waits being common.

Mad Tea Party Spin around and around to the tune of “A Very Merry Unbirthday” from Alice in Wonderland. You control how fast you spin, from moderate to very fast. Younger children often enjoy watching the world spin by. Lines are erratic, and vary in length throughout the day. This ride is not recommended for riding right after meals.

Dumbo the Flying Elephant Ride around and around, up and down, on the back of the world- famous Dumbo. The lines here are the worst in Fantasyland; while there may not be many people in line, the time for each ride, combined with the low capacity, makes the line moves very slowly. Younger children love this ride, however, and are often willing to wait in lines much longer than any parent expects them to be able to handle, all for a few moments riding on Dumbo’s back.

Matterhorn Bobsleds Ride through and around the Matterhorn mountain; this was the first of Disneyland’s mountain range of roller coasters. The inside of the mountain is realistically cold, and even has an abominable snowman or two. This is a good ride to go on with a loved one, since the seating setup allows two people to share a seat. This ride also gives a panoramic view of Disneyland from above; however, quite frequently you’re too busy wondering what will happen next to notice. There are two tracks: the left track (facing the mountain through the turnstiles) has more turns and better views of the park, while the right track has more dips and drops. Lines are not great, but look much worse than they actually are. Thirty to forty minutes waits are normal. Lines look even worse during parades, but move twice as fast because both tracks are fed from one line during parades. Anyone seven or more years old can go by themselves; anyone at least three years old can go with an adult. Riders with back, neck, or heart problems and expectant mothers are advised not to ride. Near the big ice crystals inside the mountain, a box marked “Wells Expedition” sits next to some snowshoes. This is a tribute to the late Frank Wells, president of the Walt Disney Company until his untimely death in 1993. The mountain was chosen as the tribute site because Frank Wells was a world-renowned mountaineer, having conquered the highest peak on six of the seven continents. The mountain was built to be a 1/100 scale replica of the real Matterhorn mountain, and is climbed every day by professional mountain climbers, weather permitting.

It’s a Small World Ride the Seven Seaways on the Happiest Cruise that Ever Sailed! This is one of the most famous rides in Disneyland, with a wonderfully catchy tune. Take a long (fifteen minutes) boat ride through singing, dancing Audio-Animatronic figures from around the world. It’s delightful and fun, and the lines are virtually nonexistent. If this ride isn’t virtually a walk-on, try again later. This is a great ride for large groups, since the boats hold many people at once.

Fantasyland Autopia This is a chance for everyone to drive their own car on a long, winding stretch of highway. This ride is only open during peak times; the Tomorrowland Autopia is open almost all the time. Ride restrictions are the same as for the Tomorrowland Autopia.

Fantasyland Theatre This outdoor amphitheater currently shows the Spirit of Pocahontas Stage Show. Check the daily schedule for times. There is also a walk-up window eatery next to the theatre, called Meeko’s.

Mickey’s Toontown

This is the newest land, and one of the busiest. Hit this land early or late, because the kids love this land, and at midday this place is packed. This land is the home of the cartoon characters, and many of them can be found wandering around in here. Once you enter, the “real world” goes away behind cartoon horizons. Be sure to check out the ‘streets’ around the Gag Factory; handles, boxes, doorbells, and manhole covers are all set up to provide an interactive, fun experience for kids of all ages.

Jolly Trolley Bounce from one end of Toontown to the other in these giant wind-up trolleys.

Goofy’s Bounce House Jump around inside a house that looks strangely like its owner. This is one of the few places in Disneyland with a maximum height restriction; no one over 52 inches is allowed inside.

Gadget’s Go-Coaster This is a short, fast roller coaster; it’s a great introduction to roller coasters for those too young for the Mountain Range. Be sure to look over what the coaster is made of while you’re in line; the whole view is entertaining. Riders must be at least three years old.

Chip and Dale Tree Slide and Acorn Crawl Climb into the treehouse, then go down the slide, then crawl through the yellow and brown acorns. This is another attraction with a maximum height; the acorn crawl is limited to those under 49 inches tall.

The Miss Daisy This is Donald’s boat, docked in Toon Lake; it bears a striking resemblance to its owner. Climb around, steer the boat, and slide back down.

Mickey’s House Wander through the residence of the most famous mouse in the world, then go back to his Movie Barn to meet Mickey himself. Bring a camera, because Mickey’s always here and happy to have a picture taken with you. Cast members will take the picture, if you’d like, allowing your whole group to get in the picture.

Minnie’s House Have fun wandering through Minnie’s house and seeing how Mickey’s better half lives. Design fashions for her, or see what’s in the fridge. There are many fun things for children to look at and do, so take your time.

Disneyland Railroad This is a railroad stop right outside the entrance to Toontown; this used to be the Fantasyland/Videopolis station.

Roger Rabbit’s Car Toon Spin This wildly-spinning ride is a combination of a “tracked car” ride, like Mr. Toad, and the Mad Tea Party, where you control how fast you spin. Help Roger save Jessica, but watch out for the Dip! Lines are bad (up to ninety minutes), so go early or during a parade or show. The queue is long but entertaining, and helps keep younger children amused during the wait. This ride is not recommended right after a meal.

Tomorrowland

Tomorrowland is dedicated to looking toward the future, but since the future keeps becoming the present, it’s hard to keep it up-to-date. Because of this, Tomorrowland is due for a massive makeover, rumored to be starting in January 1996. Rumored to be gone are CircleVision (replaced with “Time Keeper”), Captain EO (replaced by “Honey I Shrunk The Audience”), and the PeopleMover (replaced by “Speed Cycles”). In addition, the overall appearance of Tomorrowland will be changed to “The Future That Never Was,” as envisioned by artists earlier this century.

Even with all the planned reconstruction here, it’s still a very entertaining land, with some of the most popular rides and stores. For example, the Star Trader shop is generally considered to be second only to the Emporium on Main Street in diversity of selection, while Space Mountain and Star Tours remain two of the most popular rides. The Starcade is also one of the most up-to-date arcades around, with all sorts of new games.

Rocket Jets Ride around and around high above Tomorrowland in these small little rocket ships; this ride is very similar to Dumbo, but higher and faster. The lines are somewhat long, so riding in the morning or at night is recommended. This attraction also provides an excellent view of the park, since it’s so high off the ground.

Circle-Vision This 360-degree theater shows two different movies through the day: Wonders of China (morning and afternoon) and American Journey (afternoon and evening). It’s not recommended for small children; there are no seats, so they’re unlikely to see much with all the adults in the way. The pre-show area is a great place to rest during the heat of the day, and makes a decent meeting place during non-peak hours.

Space Mountain Ride through space in the dark, with stars and meteor showers flashing above you. This is a rollercoaster-style ride, with many sudden drops and turns, but nothing worse. Hold onto your hats, glasses, purses, and anything else that’s loose, because things have a tendency to fly away from you while you’re on this ride. Go early or late, because the lines get bad here; 60 minutes is not uncommon. Riders must be at least 40 inches tall; people with back, neck or heart problems and expectant mothers are advised not to ride.

Star Tours Take a flight simulator ride of a pleasant little trip to Endor… with a few unplanned detours. Preboarding entertainment is provided by R2-D2 and C-3PO, a repair droid, and the PA system calling for a Egroeg Sacul (George Lucas backwards) or the owner of landspeeder THX-1138 (parked in a “No-Hover” zone). Lines here tend to get very bad; 45 minutes is normal, and 60 to 90 minutes is not uncommon. Go early, late, or during a parade, but this is a must-see ride. Lines here usually are at the lowest in the late evening, especially after the first Fantasmic! show. The minimum age to ride alone is seven years old; however, the minimum with an adult is three years old. People with back, neck, or heart problems and expectant mothers are advised against going on this ride. As part of the Tomorrowland refurbishing, Star Tours is rumored to be getting a new ride film. This time, the trip is to Dagobah, and nothing will go wrong this time. Right?

PeopleMover Take a pleasant ride around Tomorrowland and see all the other rides. There is a small Tron section consisting of movies projected onto the wall you’re going through, as well as descriptions and views of virtually every other ride in Tomorrowland. Lines are short, and the ride is reasonably entertaining.

Disneyland Monorail This takes you to the Disneyland Hotel and back, complete with commentary on the sights. The ride to the hotel is uneventful, but the ride back winds through much of Tomorrowland as well as passing the Indiana Jones ride building in Adventureland. People waiting at the very front of the monorail may be able to sit up front with the driver, if they ask a Cast Member and nobody else has already claimed it.

Tomorrowland Autopia Ride around on a small track, where anyone over the height requirement can drive. Lines are moderate, and move relatively quickly. Anyone at least 52 inches tall can drive by themselves; anyone else who’s at least a year old can go with someone who is tall enough. Younger children love this ride; where else can they drive around and bump into the car in front of them? Bumping is officially not allowed and strongly discouraged, but happens frequently anyway; be forewarned if you’re subject to back or neck problems.

Submarine Voyage Cruise underwater to the North Pole, then on to Atlantis. Lines can get bad; ride early or late. You can look into the lagoon from the shore, but it isn’t the same. This ride is not recommended for those with claustrophobia.

Captain EO This is a 3D movie, starring Michael Jackson and lots of great special effects. Lines are rarely very bad, and it’s worth seeing once if only for the spectacular 3D effects. It’s due to be replaced soon with “Honey, I Shrunk the Audience,” so see it while you can. Be aware, however, that it is very loud.

Disneyland Railroad This is the last stop before the railroad returns to Main Street; the sights between Tomorrowland and Main Street include the impressive dioramas of the Grand Canyon, as well as Primeval World’s dinosaurs.

Quotes

Disneyland is going to be a place where you can’t get lost or tired

unless you want to. – Walt Disney

I don’t want the public to see the world they live in

while they’re in the Park. I want them to feel

they are in another world. – Walt Disney

Disneyland will never be completed. It will continue to grow

as long as there is imagination left in the world. – Walt Disney

I could never convince the financiers that Disneyland was feasible,

because dreams offer too little collateral. – Walt Disney

Give the people everything you can give them.

Keep the place as clean as you can keep it. Keep it friendly, you know. Make it a real fun place to be.

– Walt Disney

Disneyland is not just another amusement park.

It’s unique, and I want it kept that way. – Walt Disney

Disneyland is more than the fulfillment of one man’s dreams. It is the creation of thousands of hands and willing hearts,

an authentic wonder of the entertainment world imbued with magic. – Disneyland: Dreams, Traditions, and Transitions

Disneyland creates a thousand impressions for every guest

to cherish and remember. The dream lives on.

– Disneyland: Dreams, Traditions, and Transitions

Other Resources/Bibliography

First and foremost, Disneyland Guest Services’ recorded information is available 24 hours a day at (714) 999-4565, and it can answer most of your questions. For personal assistance, operators are available at (714) 999-4560 from 7am to 10pm daily. The address is 1313 Harbor Boulevard, P. O. Box 3232, Anaheim, CA 92803-3232. They can send maps, schedules, lists of rides unavailable due to maintenance, and lots of other information.

The Anaheim Area Visitor and Convention Bureau has information on local hotels and other places to stay. It can be reached at 800 W. Katella Avenue, Anaheim, CA 92802, or by phone at (714) 999-8999.

Some books you may want to look into are:

Birnbaum’s Disneyland: The Official Guide. This is the Disneyland guide approved by Disney. Therefore, it’s the only guide allowed to show pictures of Disneyland and its rides and events. For a good look at what you’re getting yourself into, Birnbaum’s is the only way to go. If you want information on Disneyland’s landscaping, this book most likely has it; however, its information about lines is erratic at times. (1994 edition ISBN 1-56282-804-5)

The Unofficial Guide To Disneyland, by Bob Sehlinger. This guide prides itself on its independence from Disney; however, this often translates into ignoring the Disney magic while concentrating on any flaws that can be found. He makes some very good points, but he seems to take lines as insults and treats itineraries as holy gospel. (1993 edition ISBN 0-13-953951-4)

Disneyland and Southern California with Kids, by Kim Wiley. A very good guide that covers all the necessities of bringing children to Disneyland, including what to take and where the best bathrooms are. Recommended for everyone, especially parents bringing children. This guide also covers most of the other parks in the area, at least to a small degree. (1992 edition ISBN 1-55958-296-0)

Disneyland and Beyond: the Ultimate Family Guidebook This guide isn’t quite as good as Wiley’s guide, but is still useful for parents planning to visit many area attractions. For those looking exclusively for Disneyland information, this guide is lacking; less than half the guide is devoted to Disneyland. (1992 edition ISBN 0-915233-67-3)

The Art of Walt Disney: From Mickey Mouse to the Magic Kingdoms by Christopher Finch has a pictorial section on Disneyland and Walt Disney World. Overall, this book is a nice pictorial overview of Walt’s career.

Disneyland: Inside Story by Randy Bright is a wonderful book, telling the story of how Disneyland came to be and how it became what it is today. Many pictures show the Disneyland of yesterday and today. I’ve been told it’s hard to find now, so your best bet might be to borrow it from a nearby, large library. (ISBN 0-8109-0811-5)

Disneyland: Dreams, Traditions, and Transitions by Leonard Shannon and Shelley Higuera is another excellent book about Disneyland. It is sold throughout Disneyland, and is a light pictorial and historical view of the park, with emphasis on how the park is today. This book is due to be phased-out soon, hopefully to be replaced by a newer book with even more up-to-date information.

Mouse Tales: A Behind-The-Ears Look At Disneyland by David Koenig is a very unofficial look at what actually, and allegedly, has happened at Disneyland. Many of the stories are of the type “Well, I heard that many years back someone did this and that, but we don’t do things like that anymore,” but everything is researched about as well as it could be. The section on lawsuits against Disneyland, for example, is very thoroughly researched from Orange County court records. Recommended only for people who don’t mind seeing some of the less-pleasant things that have happened at the Magic Kingdom.

Al Lutz <AlWeHo@aol.com> maintains the Disneyland Information Guide (DIG), which is a spectacular source of information on Disneyland. It is posted regularly to rec.arts.disney (soon to be

on rec.arts.disney.parks, and available via FTP from my FTP site at ftp://ftp.io.com/pub

/usr/kazrak/Disney/Disneyland/D-I-G. The rumors about Tomorrowland, among other things, come from the DIG, but have also been substantiated by a number of Company cast members.

Walt Hoffman <whoffman@ganymede.jpl.nasa.gov> maintains the Disneyland Frequently Asked Question List, which includes up-to-date information about closed rides and operating hours; the

list is available by anonymous FTP from rtfm.mit.edu in /pub/usenet/news.answers/disney-
faq/disneyland.

Tom Tanida <tanida@gso.saic.com> maintains the Walt Disney World Frequently Asked Questions List, which also contains some information about Disneyland. The list is available by anony-

mous FTP from rtfm.mit.edu in /pub/usenet/news.answers/disney-faq/disneyworld.

Legal Stuff

The following are all registered trademarks of the Walt Disney Company, Inc.: Adventureland, Audio-Animatronics, Captain EO, Disneyland, Fantasyland, Magic Kingdom, New Orleans Square, PeopleMover, Space Mountain, Walt Disney, Walt Disney World. There may be other trademarks; I will add them to the list as I become aware of them.

This guide is in no way supported or controlled by the Walt Disney Company, Inc. or any subsidiary thereof. All opinions are solely those of the author and should not be taken to be official views of the Walt Disney Company, Inc. or any subsidiary or employee thereof.

I have attempted to make this guide as accurate as possible; however, as Disneyland says on virtually all their information sheets, “Information is subject to change without notice.” The author takes no responsibility for any loss, injury, or inconvenience caused by use of the information in this guide.

This guide is Copyright \begin{matrix}c \\ ◯ \end{matrix}1994-1995 by Roger B. Jones, Jr. It may be distributed freely so long as it is not altered and the following terms are complied with. It may not be redistributed for profit without specific written consent of the author. It may not be included in a compilation of files distributed for profit without specific written consent of the author. All rights reserved.

To translate this into plain English, this guide is mine. Some of the words are Disney’s trademarks, and I don’t want to challenge that. If something goes wrong because of information in this guide, it’s not my problem, nor is it Disney’s problem. You can give the guide away for free, but if you want to make money off it, you need my permission first.

Acknowledgments

I’d like to thank the following people for their help in preparing this: My mother, for telling me of a magic place where dreams come true when I was a child, and then taking me to that place. My wife, Jennefer, for taking me back to that place several times, putting up with me writing this, and making me rewrite many parts of it that needed it. Mike and Carol, for funding these more recent expeditions and helping with the beta version of the guide. Todd McCartney, who gave me the inspiration to write this. Walt Hoffman, Steven Soria, Tom Tanida, and Laura Dolson, for helping me edit the beta version of the guide. Lisa Shoup, for commentary and editorial help above and beyond the call of duty. Rob Fish, for continual help with the guide’s layout and visual design. Al Lutz, for producing the Disneyland Information Guide, which I use for current information and rumor. The members of rec.arts.disney and rec.arts.disney.parks, for support and information. The cast members at Disneyland, for keeping the dream alive 365 days a year. And finally and most importantly, Walt Disney, the man who made the very existence of the place described here possible.